OpenStudy (anonymous):

using words how to solve the equation 6-4x =18

OpenStudy (anonymous):

subtract both sides by 6 then divide both sides by negative four. -3 is the result.

OpenStudy (anonymous):

do you kno what properties you used

jhonyy9 (jhonyy9):

6-4x=18 -4x=12 x=12/-4 x=-3